### v1

# A monolithic first–order BSSNOK formulation of the Einstein–Euler equations and its solution with path-conservative finite difference CWENO schemes

Abstract
We present a new, monolithic first–order (both in time and space) BSSNOK formulation of the coupled Einstein–Euler equations. The entire system of hyperbolic PDEs is solved in a completely unified manner via one single numerical scheme applied to both the conservative sector of the matter part and to the first–order strictly non–conservative sector of the spacetime evolution. The coupling between matter and space-time is achieved via algebraic source terms. The numerical scheme used for the solution of the new monolithic first order formulation is a path-conservative central WENO (CWENO) finite difference scheme, with suitable insertions to account for the presence of the non–conservative terms. By solving several crucial tests of numerical general relativity, including a stable neutron star, Riemann problems in relativistic matter with shock waves and the stable long-time evolution of single and binary puncture black holes up and beyond the binary merger, we show that our new CWENO scheme, introduced two decades ago for the compressible Euler equations of gas dynamics, can be successfully applied also to numerical general relativity, solving all equations at the same time with one single numerical method. In the future the new monolithic approach proposed in this paper may become an attractive alternative to traditional methods that couple central finite difference schemes with Kreiss-Oliger dissipation for the space-time part with totally different TVD schemes for the matter evolution and which are currently the state of the art in the field.
pacs:
04.25.D-, 04.25.dg,
I Introduction
Since the detection of GW150914 [1], Gravitational Waves (GWs) astronomy has experienced a fundamental boost [2], with 93 sources detected up to March 2024111https://gwosc.org/. The accurate numerical modeling of gravitational waveforms has become a key ingredient to extract signals from detector noises and to understand the properties of astrophysical sources, specifically binary systems [3, 4]. In this respect, numerical general relativity has been providing an invaluable contribution to the scientific progress (see, among the others, [5, 6, 7, 8, 9, 10, 11, 12]).
The two pillars of any successful numerical simulation are represented by an appropriate formulation of the equations that needs to be solved and by the availability of an accurate and robust numerical scheme. Concerning the formulation of the Einstein equations, the vast majority of numerical relativity codes that are currently adopted to compute gravitational waveforms is based on the celebrated (second–order in space) BSSNOK formulation [13]. In spite of a few deprecated defects, the most annoying of which is perhaps the failure in keeping a simple Kerr black hole stationary over a long timescale, this formulation has shown a spectacular success in reproducing the merger of binary systems, which was the real busillis of numerical relativity (NR). Precisely because the second–order implementation of the BSSNOK formulation has been so successful in modeling the gravitational wave signals from binary systems, the scientific research on first–order formulations has been somewhat left aside during the last decade or more. As a matter of fact, most of the analysis about first–order formulations of the Einstein equations belong to the beginning of the NR era [14, 15, 16, 17, 18, 19, 20, 21, 22], with only few exceptions in relatively recent years [23, 24, 25, 26, 27].
Concerning the numerical scheme, on the other hand, it is fair to say that, in spite of considerable efforts carried out in the last decade for adopting innovative numerical schemes in the solution of the Einstein equations, most of the existing numerical relativity codes still rely on traditional central finite differencing with Kreiss-Oliger dissipation, in combination with Runge–Kutta time integration. Of course there are exceptions to this trend. For example, there is a active field of research to extend discontinuous Galerkin (DG) schemes to full general relativity [28, 29, 30, 24, 31, 32, 33, 27]; or to develop spectral methods both for the initial value problem and for the time evolution of the Einstein equations [34, 35, 36, 37]; or to go beyond Runge–Kutta schemes for the time integration by migrating to so–called ADER schemes [38, 39, 40, 41, 42, 43]. In spite of these progresses, if we list the most popular numerical relativity codes that perform actual calculations of waveforms from binary mergers (Einstein–Toolkit [^2] http://einsteintoolkit.org/ [44], LazEv [^3] https://ccrg.rit.edu/content/software/lazev [45, 46], BAM [^4] http://data.cardiffgravity.org/bam-catalogue/ [47, 48, 49], GRChombo  [^5] https://www.grchombo.org/ [50, 51], AMReX [^6] https://amrex-codes.github.io/amrex/ [52], Nmesh [33], SACRA  [^7] https://www.aei.mpg.de/1095648/SACRA-description[53, 54], SpEC [^8] https://www.black-holes.org/code/SpEC.html [55, 35, 56], SPHINCS_BSSN [57], SENR/NRPy [58]) we observe that, apart from Nmesh and SpEC, all of them still uses finite difference schemes for the Einstein sector of the spacetime evolution.
Now, if we combine our considerations about the formulation of the Einstein equations, which rests on the second–order BSSNOK system, and the choice of a suitable numerical scheme, which rests mainly on traditional finite differencing, we see that the lack of a reliable first–order formulation represents a serious obstacle for the application of the whole class of advanced numerical schemes of Godunov type proposed in the last three decades (or more) for the solution of first–order hyperbolic PDE systems. As an additional complication, when the Einstein equations are cast in first–order form, they naturally come up with non–conservative terms and curl-type involutions, which may give the (erroneous) impression of being untractable by Godunov–type methods. On the contrary, since all the characteristic fields arising from the Einstein equations are linearly degenerate, discontinuities cannot develop, and therefore non–conservative terms in the PDE system can be taken into account in a relatively easy way. With the benefit of hindsight, we could say that insisting in finding first–order and conservative formulations of the Einstein equations has been counterproductive.
In light of the above comments, assuming that finite difference numerical schemes are likely to remain the preferred choice by the NR community for a certain amount of time onward, partly due to their simplicity and certainly for their efficiency in terms of CPU time performances, in this paper we resume a central WENO (CWENO) finite difference numerical scheme, proposed in a different context originally by [59, 60, 61], that may contribute to a substantial improvement with respect to the present standard in a few specific aspects:
  • It relies on a hyperbolic first–order but non–conservative version of the BSSNOK formulation.
  • It applies unmodified both to the spacetime (Einstein) and to the matter (Euler) evolution, achieving high order of accuracy in both space and time.
Our investigation follows some very recent progress made in [62, 63, 64], who have opened a rather innovative line of research, by introducing a new class of finite difference WENO schemes for first–order non–conservative systems. Because such schemes are quite general and not limited to NR, we address the interested reader to [65, 66, 67, 60, 61, 68, 69, 70, 71, 72], where all possible information about these new methods can be found.
At this point we would like to emphasize that, to the best knowledge of the authors, this is the first time that a fully coupled first order hyperbolic BSSNOK formulation of the Einstein-Euler equations is proposed and which is solved in a totally monolithic manner within one and the same computer code using exactly the same numerical scheme for both matter and spacetime.
The structure of the paper is as follows. In Section II we present the new first order version of the BSSNOK formulation of the coupled Einstein-Euler equations. In Section III we describe the new numerical scheme that can be used in a monolithic way for the full system of the Einstein–Euler PDEs. In Section IV we show a number of benchmark results to demonstrate the correctness of both the formulation and the numerical solver. Finally, the conclusions are summarized in Section V.
We work in a geometrized set of units, in which the speed of light and the gravitational constant are set to unity, i.e., $c=G=1$. Greek indices run from 00 to $3$, Latin indices run from $1$ to $3$ and we use the Einstein summation convention of repeated indices.
II The BSSNOK system
II.1 The standard second order version
The vast majority of numerical codes solving the Einstein–Euler equations within the $3+1$ formalism adopt the so–called BSSNOK formulation, which, due to its popularity, does not require an extended discussion and it can be briefly described as follows (for more details see  [73, 13, 74, 75]). The spatial metric $\gamma_{ij}$ of the spacelike hypersurfaces is rescaled conformally according to
$\tilde{\gamma}_{ij}=\psi^{-4}\gamma_{ij}=e^{-4\phi}\gamma_{ij}\,,$ (1)
where $\tilde{\gamma}_{ij}$ has unit determinant, such that $\psi=\gamma^{1/12}$, where $\gamma$ denotes as usual the determinant of $\gamma_{ij}$. Hence, the spatial metric $\gamma_{ij}$, with only 6 independent components, generates in fact 7 evolved quantities, namely the factor $\phi$, related to the determinant of $\gamma_{ij}$ as $\phi=\ln\psi=\frac{1}{12}\ln\gamma$, and the rescaled unit determinant metric $\tilde{\gamma}_{ij}$. Similarly, the symmetric extrinsic curvature $K_{ij}$, with only 6 independent components, generates in fact 7 evolved quantities, namely the trace $K$ of $K_{ij}$, and the rescaled trace-free tensor
$\tilde{A}_{ij}=\psi^{-4}A_{ij}=e^{-4\phi}(K_{ij}-\frac{1}{3}\gamma_{ij}K)\,.$ (2)
Overall, the (redundant) 17 evolved quantities within the BSSNOK framework obey the following system of PDEs, which is of first order in time and of mixed first and second order in space:
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{\gamma}_{ij}$ $=$ $-{2\,\alpha}\,\tilde{A}_{ij}\,,$ (3)
$(\partial_{t}-{\cal L}_{\beta})~{}\phi$ $=$ $-\frac{1}{6}\alpha\,K\,,$ (4)
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{A}_{ij}$ $=$ $e^{-4\phi}\bigg{[}-\nabla_{i}\nabla_{j}\alpha+\alpha\,R_{ij}-8\pi\alpha S_{ij}\,\bigg{]}^{TF}+\alpha\left(K\,\tilde{A}_{ij}-2\tilde{A}_{im}\tilde{A}^{m}_{j}\right)\,,$ (5)
$(\partial_{t}-{\cal L}_{\beta})~{}K$ $=$ $-\nabla_{i}\nabla^{i}\alpha+\alpha\left(\tilde{A}_{ij}\tilde{A}^{ij}+\frac{1}{3}K^{2}+4\pi(E+S)\right)\,,$ (6)
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{\Gamma}^{i}$ $=$ $\tilde{\gamma}^{jk}\partial_{j}\partial_{k}\beta^{i}+\frac{1}{3}\tilde{\gamma}^{ij}\partial_{j}\partial_{k}\beta^{k}-2\tilde{A}^{ij}\partial_{j}\alpha$
$+$ $2\alpha\left(\tilde{\Gamma}^{i}_{jk}\tilde{A}^{jk}+6\tilde{A}^{ij}\partial_{j}\phi-\frac{2}{3}\tilde{\gamma}^{ij}\partial_{j}K-8\pi e^{4\phi}M^{i}\right)$
where, according to the standard notation of the $3+1$ formalism, $\alpha$ is the lapse, $\beta^{i}$ is the shift, $R_{ij}$ is the purely spatial Ricci tensor, $S_{ij}$ is the purely spatial part of the energy momentum tensor, $S$ is its trace and the superscript $TF$ stands for the trace-free part of a tensor. Moreover, ${\cal L}_{\beta}$ denotes the Lie derivative along the shift vector. We also recall that Eq. (5) and (6) incorporate the Hamiltonian constraint, while Eq. (II.1) incorporates the momentum constraint. In terms of the BSSNOK variables these constraints can be written as
$H$ $=$ $R-\tilde{A}_{ij}\tilde{A}^{ij}+\frac{2}{3}K^{2}-16\pi E=0\,,$ (8)
$M^{i}$ $=$ $\partial_{j}\tilde{A}^{ij}+\tilde{\Gamma}^{i}_{jk}\tilde{A}^{jk}+6\tilde{A}^{ij}\partial_{j}\phi-\frac{2}{3}\tilde{\gamma}^{ij}\partial_{j}K-8\pi S^{i}=0\,,$ (9)
and these quantities will be monitored during a numerical simulation. In particular, we will check the $L_{2}$ errors of the above constraints, computed as
$L_{2}=\int_{\Omega}\,\epsilon^{2}\,\sqrt{\gamma}\,d^{3}x\,\hskip 28.45274pt\textrm{with}\hskip 28.45274pt\epsilon=H,M^{i}\,,$ (10)
where no normalization is performed with respect to the volume of integration.
II.2 The new first–order version
A promising first–order formulation of BSSNOK was proposed by [23], though it seems that traditional second order formulations are still the preferred choice in large scale simulations of astrophysical sources. Here, as already done in the first–order CCZ4 formulation of [24], we introduce 30 auxiliary variables containing first derivatives of the metric terms, namely
$A_{k}:=\partial_{k}\ln\alpha=\frac{\partial_{k}\alpha}{\alpha}\,,\qquad B_{k}^{\,\,i}:=\partial_{k}\beta^{i}\,,\qquad D_{kij}:=\frac{1}{2}\partial_{k}\tilde{\gamma}_{ij}\,.\qquad P_{k}:=\frac{\partial_{k}\psi}{\psi}=\partial_{k}\phi=\frac{1}{12}\partial_{k}\ln\gamma$ (11)
Since the auxiliary variables are gradients of primary evolution quantities, they must remain curl-free for all times if they are initially curl-free (this is obvious either via the Schwarz theorem on the symmetry of second derivatives, or alternatively because the curl of a gradient is zero). Hence, we have the following second-order ordering constraints induced by (11):
$\mathcal{A}_{lk}=\partial_{l}A_{k}-\partial_{k}A_{l}=0,\quad\mathcal{B}_{lk}^{i}=\partial_{l}B_{k}^{\,\,i}-\partial_{k}B_{l}^{\,\,i}=0,\quad\mathcal{D}_{lkij}=\partial_{l}D_{kij}-\partial_{k}D_{lij}=0,\quad\mathcal{P}_{lk}=\partial_{l}P_{k}-\partial_{k}P_{l}=0.$ (12)
Since $\tilde{A}_{ij}$ is trace-free one has $\tilde{\gamma}^{ij}\tilde{A}_{ij}=0$ and therefore another differential constraint
$\mathcal{C}_{k}=\partial_{k}\tilde{\gamma}^{ij}\tilde{A}_{ij}+\tilde{\gamma}^{ij}\partial_{k}\tilde{A}_{ij}=0$ (13)
that can be used in the first order version of BSSNOK. With the aid of (11), the system (3)–(II.1), augmented by the matter part and by the gauge conditions, can be written as a monolithic first–order BSSNOK formulation of the Einstein-Euler system, i.e.,
$\partial_{t}(\sqrt{\gamma}D)+\partial_{i}\left[\sqrt{\gamma}(\alpha v^{i}D-\beta^{i}D)\right]=0\,,$ (14)
$\partial_{t}(\sqrt{\gamma}S_{j})+\partial_{i}\left[\sqrt{\gamma}(\alpha S^{i}_{\,\,j}-\beta^{i}S_{j})\right]=\sqrt{\gamma}\left[\alpha S^{ik}D_{jik}+S_{i}B_{j}^{\,\,i}-\alpha EA_{j}\right]\,,$ (15)
$\partial_{t}(\sqrt{\gamma}E)+\partial_{i}\left[\sqrt{\gamma}(\alpha S^{i}-\beta^{i}E)\right]=\sqrt{\gamma}\left[\alpha S^{ij}e^{4\phi}\left(\tilde{A}_{ij}+\frac{1}{3}\tilde{\gamma}_{ij}K\right)-\alpha S^{j}A_{j}\right]\,,$ (16)
$\partial_{t}\tilde{\gamma}_{ij}-\beta^{k}\partial_{k}\tilde{\gamma}_{ij}=\tilde{\gamma}_{ik}B_{j}^{\,\,k}+\tilde{\gamma}_{kj}B_{i}^{\,\,k}-\frac{2}{3}\tilde{\gamma}_{ij}B_{k}^{\,\,k}-2\alpha\tilde{A}_{ij}\,,$ (17)
$\partial_{t}\phi-\beta^{k}\partial_{k}\phi=\frac{1}{6}B_{k}^{\,\,k}-\frac{1}{6}\alpha K\,,$ (18)
$\partial_{t}\tilde{A}_{ij}-\beta^{k}\partial_{k}\tilde{A}_{ij}+\alpha e^{-4\phi}\left(\partial_{(i}{A}_{j)}-\frac{1}{3}\tilde{\gamma}_{ij}\tilde{\gamma}^{mn}\partial_{(m}{A}_{n)}\right)-\alpha e^{-4\phi}\bigg{[}(R_{ij})^{TF}_{ncp}\bigg{]}=$
$\qquad\qquad\tilde{A}_{ik}B_{j}^{\,\,k}+\tilde{A}_{kj}B_{i}^{\,\,k}-\frac{2}{3}\tilde{A}_{ij}B_{k}^{\,\,k}-\alpha e^{-4\phi}\left[A_{i}A_{j}-\Gamma^{k}_{ij}A_{k}-\frac{1}{3}\tilde{\gamma}_{ij}\tilde{\gamma}^{mn}(A_{m}A_{n}-\Gamma^{k}_{mn}A_{k})\right]+$
$\qquad\qquad+\alpha e^{-4\phi}\bigg{[}(R_{ij})^{TF}_{src}\bigg{]}-8\pi\alpha e^{-4\phi}\left(S_{ij}-\frac{1}{3}e^{4\phi}\tilde{\gamma}_{ij}\,S\right)+\alpha(K\tilde{A}_{ij}-2\tilde{A}_{il}\tilde{\gamma}^{lm}\tilde{A}_{mj})\,,$ (19)
$\partial_{t}K-\beta^{k}\partial_{k}K+\alpha e^{-4\phi}\tilde{\gamma}^{ij}\partial_{(i}A_{j)}=-\alpha e^{-4\phi}\tilde{\gamma}^{ij}\left(A_{i}A_{j}-\Gamma^{k}_{ij}A_{k}\right)+\alpha\left(\tilde{A}_{ij}\tilde{A}^{ij}+\frac{1}{3}K^{2}+4\pi(E+S)\right)\,,$ (20)
$\partial_{t}\ln\alpha-\beta^{k}{\partial_{k}\ln\alpha}=-g(\alpha)\alpha(K-K_{0})\,,$ (21)
$\partial_{t}\beta^{i}-b\,s\,\beta^{k}\partial_{k}\beta^{i}=\frac{3}{4}\,s\,b^{i},$ (22)
$\partial_{t}b^{i}-b\,s\,(\beta^{k}\partial_{k}b^{i}-\beta^{k}\partial_{k}\tilde{\Gamma}^{i})=s(\partial_{t}\tilde{\Gamma}^{i}-\eta b^{i})\,,$ (23)
$\partial_{t}\tilde{\Gamma}^{i}-s\,\bigg{[}\beta^{k}\partial_{k}\tilde{\Gamma}^{i}+\tilde{\gamma}^{jk}\partial_{(j}B_{k)}^{\,\,i}+\frac{1}{3}\tilde{\gamma}^{ij}\partial_{(j}B_{k)}^{\,\,k}-\frac{4}{3}\alpha\tilde{\gamma}^{ij}\partial_{j}K\bigg{]}=$
$\qquad\qquad s\,\bigg{[}\frac{2}{3}\tilde{\Gamma}^{i}B_{k}^{\,\,k}-\tilde{\Gamma}^{k}B_{k}^{\,\,i}-2\alpha\tilde{A}^{ij}A_{j}+2\alpha\left(\tilde{\Gamma}^{i}_{jk}\tilde{A}^{jk}+6\tilde{A}^{ij}P_{j}-8\pi e^{4\phi}M^{i}\right)\bigg{]}\,,$ (24)
$\partial_{t}A_{i}-{\beta^{k}\partial_{k}A_{i}}+\alpha g(\alpha)\left(\partial_{i}K-\partial_{i}K_{0}\right)=-\alpha A_{i}\left(K-K_{0}\right)\left(g(\alpha)+\alpha g^{\prime}(\alpha)\right)+B_{i}^{\,\,k}~{}A_{k}$ (25)
$\partial_{t}B_{k}^{\,\,i}-s\left[\frac{3}{4}\partial_{k}b^{i}+b\left(\beta^{m}\partial_{m}B_{k}^{\,\,i}\right)-\mu\,\alpha^{2}\gamma^{ij}\gamma^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)\right]=s\,b\,B_{m}^{\,\,i}B_{k}^{\,\,m}\,,$ (26)
$\partial_{t}D_{kij}-{\beta^{m}\partial_{m}D_{kij}}-\frac{1}{2}{\tilde{\gamma}}_{mi}\partial_{(k}{B}_{j)}^{\,\,m}-\frac{1}{2}{\tilde{\gamma}}_{mj}\partial_{(k}{B}_{i)}^{\,\,m}+\frac{1}{3}{\tilde{\gamma}}_{ij}\partial_{(k}{B}_{m)}^{\,\,m}+\alpha\partial_{k}{\tilde{A}}_{ij}=B_{k}^{\,\,m}D_{mij}+B_{j}^{\,\,m}D_{kmi}+B_{i}^{\,\,m}D_{kmj}$
$\qquad\qquad-\frac{2}{3}B_{m}^{\,\,m}D_{kij}-\alpha A_{k}{\tilde{A}}_{ij}+\frac{1}{3}\alpha\tilde{\gamma}_{ij}\left[\tilde{\gamma}^{nm}\partial_{k}\tilde{A}_{nm}+\tilde{A}_{nm}\partial_{k}\tilde{\gamma}^{nm}\right]\,,$ (27)
$\partial_{t}P_{i}-\beta^{k}\partial_{k}P_{i}+\frac{1}{6}\alpha\partial_{i}K-\frac{1}{6}\partial_{(i}B_{k)}^{\,\,k}=P_{k}B_{i}^{\,\,k}-\frac{1}{6}\alpha KA_{i}$ (28)
We also list a few expressions and identities that are useful when writing the above PDE system in non–conservative form:
${\gamma}$ $=$ $\textnormal{det}({\gamma}_{ij})=e^{12\phi}\,,$ (29)
$\partial_{k}{\tilde{\gamma}}^{ij}$ $=$ $-2{\tilde{\gamma}}^{in}{\tilde{\gamma}}^{mj}D_{knm}\,,$ (30)
$\tilde{\Gamma}_{ij}^{k}$ $=$ ${\tilde{\gamma}}^{kl}\left(D_{ijl}+D_{jil}-D_{lij}\right)\,,$ (31)
$\Gamma_{ij}^{k}$ $=$ $\tilde{\Gamma}_{ij}^{k}+2(\delta^{k}_{i}P_{j}+\delta^{k}_{j}P_{i})-2\tilde{\gamma}^{km}P_{m}\tilde{\gamma}_{ij}\,,$ (32)
${\tilde{\Gamma}}^{i}$ $=$ ${\tilde{\gamma}}^{jk}\,{\tilde{\Gamma}}^{i}_{jk}=\tilde{\gamma}^{im}\tilde{\gamma}^{jk}\partial_{j}\tilde{\gamma}_{mk}=-\partial_{j}\tilde{\gamma}^{ij}\,,$ (33)
$\tilde{\Gamma}_{ijk}$ $=$ $\left(D_{kij}+D_{jik}-D_{ijk}\right)\,,$ (34)
$D_{kij}$ $=$ $\frac{1}{2}(\tilde{\Gamma}_{ijk}+\tilde{\Gamma}_{jik})\,,$ (35)
$\partial_{k}\tilde{\Gamma}_{ij}^{m}$ $=$ $-2{\tilde{\gamma}}^{mn}{\tilde{\gamma}}^{pl}D_{knp}\left(D_{ijl}+D_{jil}-D_{lij}\right)+{\tilde{\gamma}}^{ml}\left(\partial_{(k}{D}_{i)jl}+\partial_{(k}{D}_{j)il}-\partial_{(k}{D}_{l)ij}\right),$ (36)
$\partial_{k}\Gamma_{ij}^{m}$ $=$ $\partial_{k}\tilde{\Gamma}_{ij}^{m}+2\delta^{m}_{i}\partial_{k}P_{j}+2\delta^{m}_{j}\partial_{k}P_{i}+4\tilde{\gamma}^{mr}\tilde{\gamma}^{ns}D_{krs}P_{n}\tilde{\gamma}_{ij}-2\tilde{\gamma}_{ij}\tilde{\gamma}^{mn}\partial_{k}P_{n}-4\tilde{\gamma}^{mn}P_{n}D_{kij}\,,$ (37)
$R^{m}_{\,\,ikj}$ $=$ $\partial_{k}\Gamma^{m}_{ij}-\partial_{j}\Gamma^{m}_{ik}+\Gamma^{m}_{lk}\,\Gamma^{l}_{ij}-\Gamma^{m}_{lj}\,\Gamma^{l}_{ik},$ (38)
$R_{ij}$ $=$ $R^{k}_{\,\,ikj}=\tilde{R}_{ij}+R^{\phi}_{ij}\,,$ (39)
$\tilde{R}_{ij}$ $=$ $-\tilde{\gamma}^{lm}\partial_{l}D_{mij}+\tilde{\gamma}_{k(i}\partial_{j)}\tilde{\Gamma}^{k}+\tilde{\Gamma}^{k}\tilde{\Gamma}_{(ij)k}+\tilde{\gamma}^{lm}(2\tilde{\Gamma}^{k}_{l(i}\tilde{\Gamma}_{j)km}+\tilde{\Gamma}^{k}_{im}\,\tilde{\Gamma}_{klj})\,,$ (40)
$R^{\phi}_{ij}$ $=$ $-2\partial_{i}P_{j}-2\tilde{\gamma}_{ij}\tilde{\gamma}^{kn}\partial_{k}P_{n}+4P_{i}P_{j}-4\tilde{\gamma}_{ij}\tilde{\gamma}^{kn}P_{k}P_{n}+2\tilde{\Gamma}^{k}_{ij}P_{k}+2\tilde{\gamma}_{ij}\tilde{\gamma}^{kr}\tilde{\Gamma}^{n}_{kr}P_{n}\,,$ (41)
$R$ $=$ $\gamma^{ij}\,R_{ij}=e^{-4\phi}\tilde{\gamma}^{ij}\,R_{ij}\,,$ (42)
$\nabla_{i}\nabla_{j}\alpha$ $=$ $\alpha A_{i}A_{j}-\alpha\,\Gamma^{k}_{ij}A_{k}+\alpha\partial_{(i}{A}_{j)}\,.$ (43)
  • Eq. (21) is the gauge condition for the lapse, which can be either the harmonic one, setting $g(\alpha)=1$, or the 1+log one, setting $g(\alpha)=2/\alpha$.
  • Eqs. (22)–(II.2) all together form the well-known Gamma–driver, which constitutes the gauge condition for the shift and it can be activated or not through the parameter $s$, either 1 or 0. The extra parameter $b$ in Eq. (22) and Eq. (23), either 1 or 0, is used to switch the convection term on or off in the evolution of $\beta^{i}$ and $b^{i}$.
  • The coefficient $\mu$ in Eq. (26) can be used to modulate the insertion of the ordering constraints into the system, specifically the red terms $\gamma^{ij}\gamma^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)$, which are zero at the continuous level due to the symmetry of second derivatives, see Eq. (12).
  • The extra red term proportional to $\left[\tilde{\gamma}^{jk}\partial_{i}\tilde{A}_{jk}+\tilde{A}_{jk}\partial_{i}\tilde{\gamma}^{jk}\right]$ on the right hand side of Eq. (II.2), is added to take into account that $\tilde{A}_{jk}$ is trace free, see also [24], and are crucial to obtain a strongly hyperbolic system.
  • The terms $(R_{ij})^{TF}_{ncp}$ and $(R_{ij})^{TF}_{src}$ in Eq. (II.2) contain the non–conservative products and the purely algebraic factors, respectively, that can be extracted from Eq. (39) when all the partial derivatives of the Christoffel symbols are expanded.
The definition of the matter quantities $D$, $E$, $S_{i}$, $S_{ij}$ for a non–dissipative fluid in Eq. (14)–(16) is pretty standard and it can be found, for instance, in [75]. Throughout this paper we are assuming a simple ideal gas equation of state $p=\rho\epsilon(\gamma-1)$, where $\epsilon$ is the specific internal energy, while $\gamma$ is the adiabatic index. Of course, present day NR simulations adopt much more realistic equations of state, especially for neutron star binaries (see, among the others, [76, 77, 78]), but this is not the focus of this work, and we have therefore set up the simplest physical conditions. For the sake of completeness we recall the definitions of the conservative variables and of the stress tensor in terms of the primitive (physical) variables:
$D=\rho W,\quad S_{i}=\rho hW^{2}v_{i},\quad E=\rho hW^{2}-p,\quad S_{ij}=\rho hW^{2}v_{i}v_{j}+p\,\gamma_{ij},\quad h=1+\epsilon+\frac{p}{\rho},\quad W=\frac{1}{\sqrt{1-v^{i}v_{i}}},$ (44)
with $h$ the specific enthalpy, $\rho$ the rest mass density, $v_{i}$ the fluid velocity and $W$ the Lorentz factor. For an efficient and robust conversion from conservative to primitive variables, even including vacuum, see [27]. The new first order BSSNOK system proposed above is strongly hyperbolic and in Appendix A we provide the list of eigenvalues and eigenvectors for a reduced subset of the full system.
III The numerical method
III.1 Central WENO reconstruction
Central WENO (CWENO) schemes were proposed more than 20 years ago in a series of papers by [59, 60, 61]. Here we provide a step-by-step presentation of this approach, showing how it can be practically implemented within the framework of conservative WENO finite difference methods [65, 66, 67]. We consider for simplicity the $x$-direction, while the extension to a 3D implementation can be easily obtained after repeating the same reconstruction also in the $y$ and $z$ directions. We assume the discrete solution of Eq. (59) below to be given by the point values $\mathbf{u}_{i,j,k}=\mathbf{u}(x_{i},y_{j},z_{k})$, with the points located in the centers of the logically equidistant control volumes, i.e. $x_{i}=\frac{1}{2}(x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}})$ and $\Delta x=x_{i+\frac{1}{2}}-x_{i-\frac{1}{2}}$. More precisely, we have
$x_{i}=x_{L}+\frac{1}{2}\Delta x+(i-1)\Delta x,\qquad\Delta x=\frac{x_{R}-x_{L}}{\textnormal{IMAX}}$ (45)
with $x_{L}$ and $x_{R}$ the left and right boundaries of the computational domain and IMAX the number of cells in the $x$ direction. Similar definitions apply to $y_{j}$ and $z_{k}$.
  1. 1.
    In each direction, there is a total stencil $\mathcal{S}_{opt}$ given by $\mathcal{S}_{opt}=\left\{i-r,\cdots,i-1,i,i+1,\cdots,i+r\right\}$ with $r=N/2$ the half stencil size and $N$ the degree of the polynomial to be reconstructed on such a stencil. In addition, there are a number of sub-stencils, which we list below according to the value of $N$:
    • For $N=2$
      $\mathcal{S}_{L}=\left\{i-1,i\right\},\hskip 179.25244pt\mathcal{S}_{R}=\left\{i,i+1\right\}\,.\hskip 44.10185pt$ (46)
    • For $N>2$
      $\mathcal{S}_{L}=\left\{i-2,i-1,i\right\},\hskip 28.45274pt\mathcal{S}_{C}=\left\{i-1,i,i+1\right\},\hskip 28.45274pt\mathcal{S}_{R}=\left\{i,i+1,i+2\right\}\,.$ (47)
  2. 2.
    A generic polynomial of degree $M$ on any stencil $k$ is expressed in terms of a set of polynomial basis functions $\psi_{m}(x)$ and associated degrees of freedom $\hat{u}_{k}^{m}$ as follows:
    $P^{M}_{k}(x)=\sum\limits_{m=0}^{M}\psi_{m}(x)\,\hat{u}^{m}_{k}\,,\qquad\textrm{with}\qquad k\in\left\{opt,L,C,R\right\}\,.$ (48)
    Throughout this paper we will use the simple Taylor monomials $\psi_{m}(x)=x^{m}$ as basis functions. The polynomials $P^{M}(x)$ are obtained by reconstruction, i.e. by interpreting the point values $\mathbf{u}_{i,j,k}$ of the finite difference scheme as cell averages, see [66]. The associated reconstruction equations (which are not interpolation equations, see [66, 67]) for a generic polynomial of degree $M$ computed on an associated stencil $\mathcal{S}_{k}$ read
[…]
    In practice, the polynomials to be reconstructed, according to the value of $N$, are chosen as:
    • For $N=2$                   $P^{2}_{opt}(x)$ ,                   $P^{1}_{L}(x)$ ,                   $P^{1}_{R}(x)$
    • For $N>2$                   $P^{N}_{opt}(x)$ ,                   $P^{2}_{L}(x)$ ,                   $P^{2}_{C}(x)$ ,                   $P^{2}_{R}(x)$
  3. 3.
    After that, we assume that the polynomial $P^{M}_{opt}(x)$, with $M\in\left\{2,N\right\}$ is split in terms of the poynomials over the sub-stencils, plus one additional (yet unknown) $P^{M}_{0}(x)$ polynomial, i.e.
    • For $N=2$
      $P^{2}_{opt}(x)=\lambda_{0}P^{2}_{0}(x)+\lambda_{L}P^{1}_{L}(x)+\lambda_{R}P^{1}_{R}(x)\,.$ (50)
    • For $N>2$
      $P^{N}_{opt}(x)=\lambda_{0}P^{N}_{0}(x)+\lambda_{L}P^{2}_{L}(x)+\lambda_{C}P^{2}_{C}(x)+\lambda_{R}P^{2}_{R}(x)\,.$ (51)
    The linear weights $\lambda_{0}$, $\lambda_{C}$, $\lambda_{L}$ and $\lambda_{R}$ in a CWENO scheme can be chosen arbitrarily. Throughout this paper we typically set $\lambda_{0}=10^{8}$, $\lambda_{C}=10^{4}$, and $\lambda_{L}=\lambda_{R}=1$, following [79], though different choices have also been considered, depending on the test, as specified later.
  4. 4.
    Consequently, from (50) and (51) we now compute the auxiliary polynomial as
    • For $N=2$
      $P^{2}_{0}(x)=\frac{1}{\lambda_{0}}\left(P^{2}_{opt}(x)-\lambda_{L}P^{1}_{L}(x)-\lambda_{R}P^{1}_{R}(x)\right)\,.$ (52)
    • For $N>2$
    $\hskip 28.45274ptP^{N}_{0}(x)=\frac{1}{\lambda_{0}}\left(P^{N}_{opt}(x)-\lambda_{L}P^{2}_{L}(x)-\lambda_{C}P^{2}_{C}(x)-\lambda_{R}P^{2}_{R}(x)\right)\,.$ (53)
  5. 5.
    Finally, a non-linear WENO reconstruction is performed in terms of the polynomial $P^{N}_{0}(x)$ and of the lower degree polynomials built over the existing sub-stencils. In particular, we compute for all polynomials of degree $M(k)$ the nonlinear WENO oscillation indicators as follows
[…]
    Tab. 1 shows the main values of the degree polynomials associated to their stencils. The nonlinear WENO weights ${\omega}_{k}$ are then computed as usual as
    ${\omega}_{k}=\frac{\tilde{\omega}_{k}}{\sum_{m}\tilde{\omega}_{m}},\qquad\textnormal{ with }\qquad\tilde{\omega}_{k}=\frac{\lambda_{k}}{\left(\sigma_{k}+\epsilon\right)^{r}}\,,$ (55)
    where we typically set $r=4$ and $\epsilon=10^{-7}$. The final CWENO reconstruction polynomial then reads
    $w_{h}(x)=\sum_{k}\omega_{k}P^{M(k)}_{k}(x).$ (56)
Table 1: Main parameters of the CWENO polynomials.
scheme order $N$ $M(0)$ $M(L)$ $M(C)$ $M(R)$
3 2 2 1 1
5 4 4 2 2 2
7 6 6 2 2 2
9 8 8 2 2 2
The reconstructed states on the left and right interface of each grid point are given component-wise and in a dimension-by-dimension fashion as
$w^{\mp}_{i\pm\frac{1}{2},j,k}=w_{h}\left(x_{i\pm\frac{1}{2}}\right)$ (57)
so that the corresponding spatial derivative can be obtained as
$\partial_{x}\boldsymbol{u}_{i,j,k}=\frac{\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{+}_{i-\frac{1}{2},j,k}}{\Delta x},$ (58)
which is a very interesting and peculiar feature of conservative WENO schemes, see [66, 67]. The same can be done also for the discrete derivatives in $y$ and $z$ direction so that we finally get the discrete spatial gradient $\nabla\boldsymbol{u}_{i,j,k}=\left(\partial_{x}\boldsymbol{u}_{i,j,k},\partial_{y}\boldsymbol{u}_{i,j,k},\partial_{z}\boldsymbol{u}_{i,j,k}\right)^{T}$ to be used in the scheme (60) shown later. The CWENO reconstruction presented above for the point values of the states $\boldsymbol{u}_{i,j,k}$ is also applied to the point values of the fluxes, i.e. for $\boldsymbol{f}^{x}_{i,j,k}$ in the $x$ direction, $\boldsymbol{f}^{y}_{i,j,k}$ in the $y$ direction and $\boldsymbol{f}^{z}_{i,j,k}$ in the $z$ direction.
III.2 CWENO finite difference discretization of the Einstein-Euler system
The full matter + Einstein equations (14)–(II.2) form a first–order PDE system in which the matter sector enters as a purely conservative sub–system, while the Einstein sector is purely non–conservative. On the overall it can be written as
$\frac{\partial\boldsymbol{u}}{\partial t}+\frac{\partial\boldsymbol{f}^{i}}{\partial x_{i}}+\boldsymbol{B}^{i}(\boldsymbol{u})\frac{\partial\boldsymbol{u}}{\partial x_{i}}=\mathbf{S}(\boldsymbol{u}),\qquad\textnormal{or, equivalently,}\qquad\frac{\partial\boldsymbol{u}}{\partial t}+\nabla\cdot\boldsymbol{F}(\boldsymbol{u})+\boldsymbol{B}(\boldsymbol{u})\cdot\nabla\boldsymbol{u}=\mathbf{S}(\boldsymbol{u})\,,$ (59)
where $\boldsymbol{u}$ is the state vector, composed of 63 dynamical variables [^9] More specifically, 5 for the matter part, 11 for the lapse, the shift vector, the metric components and the scalar $\phi$, 7 for $\tilde{A}_{ij}$ and the scalar $K$, 3 for $A_{i}$, 9 for $B_{i}^{\,\,j}$, 18 for $D_{ijk}$, 3 for $P_{i}$, 1 for $K_{0}$, 3 for $b^{i}$ and 3 for $\tilde{\Gamma}^{i}$., while $\boldsymbol{F}=\left(\boldsymbol{f}^{x},\boldsymbol{f}^{y},\boldsymbol{f}^{z}\right)^{T}$ is the flux vector. A finite difference WENO scheme for the vacuum Einstein equations, namely with $T^{\mu\nu}=0$ and which can successfully cope with the presence of non–conservative terms, has been recently proposed by [64]. Here we follow a similar approach, with a few modifications. Still in semi-discrete form, a path-conservative finite difference scheme for the discretization of (59), reads
$\frac{d\boldsymbol{u}_{i,j,k}}{dt}=$ $-\frac{\boldsymbol{f}^{x}_{i+\frac{1}{2},j,k}-\boldsymbol{f}^{x}_{i-\frac{1}{2},j,k}}{\Delta x}-\frac{\boldsymbol{f}^{y}_{i,j+\frac{1}{2},k}-\boldsymbol{f}^{y}_{i,j-\frac{1}{2},k}}{\Delta y}-\frac{\boldsymbol{f}^{z}_{i,j,k+\frac{1}{2}}-\boldsymbol{f}^{z}_{i,j,k-\frac{1}{2}}}{\Delta z}-\boldsymbol{B}(\boldsymbol{u}_{i,j,k})\cdot\nabla\boldsymbol{u}_{i,j,k}+$ (60)
$-\frac{\mathbf{D}^{x}_{i+\frac{1}{2},j,k}+\mathbf{D}^{x}_{i-\frac{1}{2},j,k}}{\Delta x}-\frac{\mathbf{D}^{y}_{i,j+\frac{1}{2},k}+\mathbf{D}^{y}_{i,j-\frac{1}{2},k}}{\Delta y}-\frac{\mathbf{D}^{z}_{i,j,k+\frac{1}{2}}+\mathbf{D}^{z}_{i,j,k-\frac{1}{2}}}{\Delta z}+\mathbf{S}(\boldsymbol{u}_{i,j,k}).$
The calculation of the discrete spatial gradient $\nabla\boldsymbol{u}_{i,j,k}=\left(\partial_{x}\boldsymbol{u}_{i,j,k},\partial_{y}\boldsymbol{u}_{i,j,k},\partial_{z}\boldsymbol{u}_{i,j,k}\right)^{T}$ is obtained through the CWENO reconstruction, as described in Sect. III.1. The calculation of the numerical fluxes $\boldsymbol{f}^{x}_{i\pm\frac{1}{2},j,k}$, $\boldsymbol{f}^{y}_{i,j\pm\frac{1}{2},k}$ and $\boldsymbol{f}^{z}_{i,j,k\pm\frac{1}{2}}$ is obtained through an approximate Riemann solver, for which we provide two possibilities
  • a Rusanov-type (local Lax-Friedrichs) Riemann solver (written here along the $x$-direction), i.e.,
    $\boldsymbol{f}^{x}_{i+\frac{1}{2},j,k}=\frac{1}{2}\left(\boldsymbol{f}^{x,-}_{i+\frac{1}{2},j,k}+\boldsymbol{f}^{x,+}_{i+\frac{1}{2},j,k}\right)-\frac{1}{2}\lambda^{\max}_{i+\frac{1}{2},j,k}\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)\,,$ (61)
  • an HLL Riemann solver (written here along the $x$-direction), i.e.,
    $\boldsymbol{f}^{x}_{i+\frac{1}{2},j,k}=\frac{\lambda^{+}_{i+\frac{1}{2},j,k}\boldsymbol{f}^{x,-}_{i+\frac{1}{2},j,k}-\lambda^{-}_{i+\frac{1}{2},j,k}\boldsymbol{f}^{x,+}_{i+\frac{1}{2},j,k}+\lambda^{-}_{i+\frac{1}{2}}\lambda^{+}_{i+\frac{1}{2},j,k}\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)}{\lambda^{+}_{i+\frac{1}{2},j,k}-\lambda^{-}_{i+\frac{1}{2},j,k}}\,,$ (62)
where
$\lambda^{-}_{i+\frac{1}{2},j,k}$ $=$ $\textnormal{min}\left(0.0,\textnormal{min}\left(\tilde{\lambda}(\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}),\tilde{\lambda}(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k})\right)\right)$ (63)
$\lambda^{+}_{i+\frac{1}{2},j,k}$ $=$ $\textnormal{max}\left(0.0,\textnormal{max}\left(\tilde{\lambda}(\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}),\tilde{\lambda}(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k})\right)\right)$ (64)
$\lambda^{\max}_{i+\frac{1}{2},j,k}$ $=$ $\textnormal{max}\left(|\tilde{\lambda}(\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k})|,|\tilde{\lambda}(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k})|\right)$ (65)
are obtained from the eigenvalues $\tilde{\lambda}$ of the system matrix $\boldsymbol{A}^{x}(\boldsymbol{u})=\partial\boldsymbol{f}^{x}/\partial\boldsymbol{u}+\boldsymbol{B}^{x}(\boldsymbol{u})$ computed at the interface as
$\tilde{\boldsymbol{A}}^{x}_{i+\frac{1}{2},j,k}=\boldsymbol{A}^{x}(\bar{\boldsymbol{w}})\,,\qquad\bar{\boldsymbol{w}}=\frac{1}{2}\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}+\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)\,.$ (66)
Note that one only needs the eigenvalues, not the matrix itself. Eq. (61) and (62) should be regarded as four–state flux formulas, where each of the four entries $\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}$, $\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}$, $\boldsymbol{f}^{x,-}_{i+\frac{1}{2},j,k}$ and $\boldsymbol{f}^{x,+}_{i+\frac{1}{2},j,k}$ are reconstructed via the CWENO strategy, as described in Sect. III.1. The whole procedure must of course be repeated for the calculation of $\boldsymbol{f}^{y}_{i,j\pm\frac{1}{2},k}$ and $\boldsymbol{f}^{z}_{i,j,k\pm\frac{1}{2}}$. According to [72, 64] the three terms highlighted in red in Eq. (60) are equivalent to the jump terms in path-conservative finite volume schemes [80, 81, 82, 83] and they are given as follows:
  • Rusanov–type dissipation:
    $\mathbf{D}^{x}_{i+\frac{1}{2},j,k}$ $=$ $\frac{1}{2}\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}\cdot\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)\,,$ (67)
  • HLL–type dissipation:
    $\mathbf{D}^{x}_{i+\frac{1}{2},j,k}$ $=$ $-\frac{\lambda^{-}_{i+\frac{1}{2}}}{\lambda^{+}_{i+\frac{1}{2}}-\lambda^{-}_{i+\frac{1}{2}}}\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}\cdot\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)\,,$
    $\mathbf{D}^{x}_{i-\frac{1}{2},j,k}$ $=$ $\phantom{-}\frac{\lambda^{+}_{i-\frac{1}{2}}}{\lambda^{+}_{i-\frac{1}{2}}-\lambda^{-}_{i-\frac{1}{2}}}\tilde{\boldsymbol{B}}^{x}_{i-\frac{1}{2},j,k}\cdot\left(\boldsymbol{w}^{+}_{i-\frac{1}{2},j,k}-\boldsymbol{w}^{-}_{i-\frac{1}{2},j,k}\right).$ (69)
The computation of the terms $\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}$ in (67) and (III.2) is performed through a simple midpoint rule as
$\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}=\boldsymbol{B}^{x}(\bar{\boldsymbol{w}})\,,\qquad\bar{\boldsymbol{w}}=\frac{1}{2}\left(\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}+\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}\right)\,.$ (70)
At first glance, one may not have expected such path-conservative jump terms in the numerical scheme and in fact they do not appear in a traditional central finite difference scheme. However, as shown by [64], in the context of WENO finite difference schemes they have the practical effect of increasing the order of accuracy by one with very little extra cost, and furthermore also substantially increase the robustness of the final numerical method and should therefore be retained.
Similar expressions for the path-conservative jump terms must also be produced for the $y$ and $z$ directions, respectively.
According to [66] the WENO approximation of the discrete gradient at the cell-center can be written via the divided differences of the boundary-reconstructed states as
$\nabla\boldsymbol{u}_{i,j,k}=\left(\begin{array}[]{c}(\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}-\boldsymbol{w}^{+}_{i-\frac{1}{2},j,k})/\Delta x\\ (\boldsymbol{w}^{-}_{i,j+\frac{1}{2},k}-\boldsymbol{w}^{+}_{i,j-\frac{1}{2},k})/\Delta y\\ (\boldsymbol{w}^{-}_{i,j,k+\frac{1}{2}}-\boldsymbol{w}^{+}_{i,j,k-\frac{1}{2}})/\Delta z\end{array}\right).$ (71)
This concludes the description of the spatial discretization used in this paper.
Finally, the nonlinear ODE system resulting from (60) is discretized in time via high order classical or TVD Runge–Kutta schemes, see [65, 66, 84].
Grid stretching.
The choice of coordinates is largely determined by the initial conditions of the problem to be solved. Most of the 3D simulations of NR adopt initial conditions that are conformally flat, and for this reason the corresponding coordinates are called Cartesian. Our simple finite difference code does not yet incorporate Adaptive Mesh Refinement (AMR), but it implements a coordinate transformation from the logical Cartesian grid to the physical domain, with the aim of producing a high resolution (but uniform) grid in the inner part, while producing a stretched non-uniform grid in the outer part, with an outer border much farther away. In practice, along each direction we perform a mapping from the logical coordinate $\xi_{i}$ to the physical coordinate $x_{i}$ simply as
$x_{i}(\xi_{i})=\left\{\begin{array}[]{llll}\xi_{i}&{\rm if}&|\xi_{i}|\leq\xi_{i}^{c}\,,\\ \frac{\xi_{i}}{|\xi_{i}|}\left(a|\xi_{i}|^{3}+b|\xi_{i}|^{2}+c|\xi_{i}|+d\right)&{\rm if}&|\xi_{i}|>\xi_{i}^{c}\,,\end{array}\right.$ (72)
where $\xi_{i}^{c}$ sets the border of the inner uniform grid, while the coefficients $a$, $b$, $c$ and $d$ are selected in such a way to obtain a smooth matching of the cubic with the inner linear branch. The same mapping is applied along all coordinate directions.
Well–balancing.
For stationary (or quasi stationary) problems, our numerical scheme can easily be made well–balanced, by simply subtracting a discrete version of the equilibrium solution from the discretized time-dependent PDE system. In this way, an initial equilibrium can be maintained stationary up to machine precision virtually for ever. The interested reader is referred to [85, 86, 27], where he/she can find all the practical details for the simple and straightforward implementation of a well–balanced scheme in a PDE system of mixed conservative///non–conservative form. For more theoretical background on well-balanced numerical methods for hyperbolic PDE, see [87, 88, 89, 90, 91, 80, 81, 82] and references therein.
IV Numerical tests
In this Section we present a large set of canonical tests for numerical general relativity, obtained with our new high order path-conservative finite difference CWENO schemes applied to the novel first order BSSNOK formulation of the coupled Einstein-Euler system. There are of course plenty of such tests that have been proposed over the year, some of which already quite standardized [92, 93]. Due to lack of space, we had to perform a selection, as outlined below. For all computational tests shown in this section we have used the classical fourth order Runge-Kutta scheme to integrate the semi-discrete CWENO method in time. Moreover, in several of the tests below we monitor the Einstein constraints of Eq. (8)–(9), referring to them for simplicity simply as $H$, $M1$, $M2$, $M3$.
IV.1 Linearized gravitational wave
We start with a simple wave perturbation of the flat Minkowski spacetime [92] for which the metric is given by
$ds^{2}=-dt^{2}+dx^{2}+(1+b)\,dy^{2}+(1-b)\,dz^{2},\quad\textnormal{with}\quad b=\epsilon\sin\left(2\pi(x-t)\right)\,.$ (73)
All the metric terms are directly deduced from (73), where, for instance, the extrinsic curvature follows from $K_{ij}=\partial_{t}\gamma_{ij}/(2\alpha)$. Moreover, since we choose $\epsilon$ to be $\epsilon=10^{-8}$, the overall dynamics is linear and the terms depending on $\epsilon^{2}$ can be neglected. We use the harmonic gauge condition, while the gamma–driver can be turned off, i.e. $s=0$.
Figure 1: Linearized gravitational wave test solved with different version of the CWENO scheme. Left panel: $\tilde{A}_{zz}$ component of the extrinsic curvature at the final time, compared to the exact solution. Right panel: time evolution of the Einstein constraints.
Matter is absent in this test. The computational domain is given by the rectangle $\Omega=[-0.5,0.5]\times[-0.05,0.05]$, which is discretized using $32\times 4$ grid-points, and adopting periodic boundary conditions in both directions.
In Figure 1 we report the results of the calculation comparing the performance of our CWENO schemes at different orders. In the left panel we show the profiles of $\tilde{A}_{zz}$ at $t=1000$, compared to the exact solution. Apart from the CWENO5 scheme, which shows an appreciable deviation with respect to the solid line, the behavior of the solutions obtained with CWENO7 and CWENO9 is very accurate. In the right panel we monitor the evolution of the Einstein constraints, which remain close to machine precision all along the simulation.
IV.2 The robust stability test
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: Robust stability test case with a random initial perturbation of amplitude $10^{-7}/\rho^{2}$ in all quantities on a sequence of successively refined meshes on the unit square in 2D. The gamma–driver shift condition, $1+\log$ slicing and CWENO5 scheme have been used. Top left: $10\times 10$ elements. Top right: $20\times 20$ elements. Bottom left: $40\times 40$ elements. Bottom right: $80\times 80$ elements.
The robust stability test introduced in [92] is used to highlight potential unstable and exponentially growing modes in the solution, and it is therefore a procedure to verify the hyperbolicity of the PDE system in a pragmatic and empirical way. It is performed on the two dimensional domain $\Omega=[-0.5;0.5]\times[-0.5;0.5]$ by perturbing a flat Minkowski spacetime with a random perturbation with amplitude ±$\pm 10^{-7}/\varrho^{2}$. The parameter $\varrho$, which is used to scale the perturbation, affects also the resolution of the grid, which is composed of $10\varrho\times 10\varrho$ gridpoints. As customary for this test, the gamma–driver shift condition is activated.
Fig. 2 shows the results of our calculations, by reporting the evolution of the Einstein constraints in a battery of tests where the CWENO5 version of the scheme is adopted. These numbers and the corresponding results are certainly satisfactory, even if not of the same quality as those reported in Fig. 4 of [27], where a DG scheme was used in combination with the Z4 formulation of the Einstein equations  [94, 22].
IV.3 The gauge wave
Figure 3: Solution of the gauge wave test at $t=1000$ with $A=0.01$ using the CWENO7 scheme. Left panel: Evolution of the Einstein constraints. Right panel: profile of the metric term $\tilde{\gamma}_{xx}$ at the final time compared to the initial condition.
The so–called gauge wave test, taken from [92], is traditionally a challenging test for the BSSNOK formulation of the Einstein equations (see [95, 23]). We found that this is still true even with our first–order BSSNOK implementation. We recall that, on the contrary, rather successful performances with this test were reported by [24] in their first–order reformulation of the CCZ4 system and more recently by [27] with an undamped first–order version of the pure Z4 formulation. In this test the metric is obtained from a simple coordinate transformation performed in the Minkowski spacetime and it is given by
$ds^{2}=-H(x,t)\,dt^{2}+H(x,t)\,dx^{2}+dy^{2}+dz^{2},\quad\text{where}\quad H(x,t)=1-A\,\sin\left(2\pi(x-t)\right)\,,$ (74)
which is in fact equivalent to a sinusoidal gauge wave of amplitude $A$ propagating along the $x$-axis. A harmonic gauge condition is required along with periodic boundary conditions.
We have first run this test with a small wave amplitude $A=0.01$ over a rectangular domain of size $\Omega=[-0.5,0.5]\times[-0.02,0.02]$ using a CWENO7 scheme. The grid is composed of $256\times 4$ elements, uniformly distributed, and the final time is $t=1000$. As expected, the gauge wave test manifests several pathologies when evolved within the BSSNOK formulation over a long timescale. We have confirmed this behaviour, which is clearly in-printed in the exponential growth of the Einstein constraints, as reported in the left panel of Figure 3. In the right panel we can appreciate the mismatch of the final solution at $t=1000$ with respect to the initial conditions, taking the metric function $\tilde{\gamma}_{xx}$ as a representative quantity.
In spite of these deficiencies, it is still possible to perform a convergence analysis upon this solution, provided the final time is short enough in such a way that the pathologies highlighted above do not have the time to spoil the numerical properties of the scheme. Tab. 2 contains the result of this analysis for a gauge wave with $A=0.1$ at the final time $t=1$ and it essentially confirms the nominal orders of convergence of the scheme.
Table 2: Numerical convergence results for the gauge wave with $A=0.1$ at a final time of $t=1$. The $L^{2}$ errors and the corresponding observed convergence order are reported for the variables $K$, $\phi$ and $\alpha$.
$N$ ${L^{2}}$ error $K$ $\mathcal{O}(K)$ ${L^{2}}$ error $\phi$ $\mathcal{O}(\phi)$ ${L^{2}}$ error $\alpha$ $\mathcal{O}(\alpha)$
CWENO3
$32$ 3.984E-04 7.963E-06 4.778E-05
$64$ 5.111E-05 3.0 1.007E-06 3.0 6.039E-06 3.0
$128$ 6.418E-06 3.0 1.261E-07 3.0 7.566E-07 3.0
$256$ 9.042E-07 2.8 1.919E-08 2.8 1.151E-07 2.7
CWENO5
$32$ 1.046E-05 1.291E-07 7.746E-07
$64$ 3.380E-07 5.0 4.135E-09 5.0 2.481E-08 5.0
$128$ 1.066E-08 5.0 1.302E-10 5.0 7.809E-10 5.0
$256$ 3.341E-10 5.0 4.079E-12 5.0 2.448E-11 5.0
CWENO7
$32$ 3.681E-06 3.591E-08 2.155E-07
$64$ 5.294E-07 6.7 5.076E-09 6.8 3.046E-08 6.8
$128$ 4.536E-09 6.9 4.330E-11 6.9 2.598E-10 6.9
$256$ 4.475E-11 6.7 4.525E-13 6.6 2.715E-12 6.6
IV.4 Special relativistic Riemann Problems in the Cowling approximation
As a first test involving the matter part of the PDE system, we consider a few canonical Riemann problems in the Cowling approximation [96], namely after neglecting the evolution of the spacetime, which is assumed to be flat Minkowski. We have selected a sample of three Riemann problems, whose initial conditions are reported in Table 3. Each of them has been solved using a CWENO3 scheme, a resolution of $512$ points and WENO parameters of Eq. (51)–(55), that, contrary to the standard values declared in Sect. III.1, are given by $\lambda_{0}=10^{8}$, $\lambda_{C}=10^{4}$, $\lambda_{L}=1$, $\lambda_{R}=1$, $r=10$, $\epsilon=10^{-14}$.
The corresponding solutions are shown in Fig. 46, and they are schematically described as
  • Riemann Problem 1, already considered by [97], produces a rarefaction wave propagating to the left, a contact discontinuity and a second rarefaction wave propagating to the right.
  • Riemann Problem 2, already considered by [98], produces a rarefaction wave propagating to the left, a contact discontinuity and a shock wave propagating to the right.
  • Riemann Problem 3, already considered by [99], produces a shock wave propagating to the left, a contact discontinuity and a second shock wave propagating to the right.
All together, these calculation prove the ability of the numerical scheme in treating strong discontinuities in the relativistic regime. At this point we would like to emphasize again that the employed conservative CWENO finite difference scheme is exactly the same as used previously for the evolution of vacuum spacetimes.
Table 3: Initial left (L) and right (R) states of the Riemann problems. The last column reports the final time $t_{f}$ of the simulation.
Problem $\rho_{L}$ $v_{L}$ $p_{L}$ $\rho_{R}$ $v_{R}$ $p_{R}$ $\gamma$ $t_{f}$
1 1 -0.6 10 10 0.5 20 5/3 0.4
2 $10^{-3}$ 0 1 $10^{-3}$ 0 $10^{-5}$ 5/3 0.4
3 1 0.9 1 1 0 10 4/3 0.4
Figure 4: Solution of Riemann Problem 1 at time $t=0.4$.
Figure 5: Solution of Riemann Problem 2 at time $t=0.4$.
Figure 6: Solution of Riemann Problem 3 at time $t=0.4$.
IV.5 Michel accretion
Again in the Cowling approximation, we can consider the motion of a gas that is accreting onto a non rotating black hole through a spherically symmetric stationary flow. A detailed presentation of the solution can be found in [75, 100]. We just note here that in Kerr–Schild coordinates the computation of the quantities at the critical radius follows the same relations valid in Schwarzschild coordinates. For example, using as free parameters the critical radius $r_{c}$ and the critical density $\rho_{c}$, the $r-$component of the four velocity at the critical radius is still given by $|u^{r}_{c}|=1/\sqrt{2r_{c}}$, just like in Schwarzschild coordinates. What changes, though, is the value of $u^{t}$, which follows from the normalization condition $u^{\mu}u_{\mu}=-1$ and is given by
$u^{t}=\frac{-zu^{r}-\sqrt{(u^{r})^{2}-z+1}}{z-1}\hskip 28.45274pt{\textrm{with}}\,\,\,\,\,u^{r}<0\,,\,\,\,\,\,z=2/r\,.$ (75)
We have solved this test over a computational domain $(r,\theta)\in[0.5;10]\times[0+\epsilon;\pi-\epsilon]$, with $\epsilon=0.05$ and covered by a $128\times 16$ uniform grid. The relevant parameters are chosen as $r_{c}=5$ and $\rho_{c}=1.006\times 10^{-7}$, while an ideal gas with adiabatic index $\gamma=5/3$ is assumed. In addition, the initial condition has been perturbed in the rest mass density using a Gaussian profile peaked at the critical radius, with a perturbation amplitude given by $\delta\rho=10^{-1}\rho_{c}$. The well–balancing property described in Sect. III.2 and proposed in [27] has been activated, in combination with the CWENO5 version of the scheme. In spite of the perturbation, and by virtue of the well–balancing, the solution relaxes back to perfect equilibrium. In Fig. 7 we have reported the final profiles of the mass density and of the radial velocity at $t=1000$.
Refer to caption Refer to caption
Figure 7: Spherical accretion of matter onto a non-rotating black hole using Kerr-Schild coordinates. The initial condition is compared to the solution at time $t=1000\,M$, both for the rest mass density (left panel) and for the radial velocity (right panel). A CWENO5 scheme with Well Balancing is adopted.
IV.6 The equilibrium TOV star
Another fundamental test in numerical general relativity consists of keeping an equilibrium star model stationary over long timescales, providing the first case in our sample of tests where the full Einstein+Euler equations are addressed. Hence, we have first solved the standard Tolman–Oppenheimer–Volkoff (TOV) system [101, 102, 75, 103] for a simple polytropic gas obeying $p=K\,\rho^{\gamma}$.
After choosing the parameters as in [104], namely a central rest mass density $\rho_{c}=1.28\times 10^{-3}$, $K=100$ and $\gamma=2$, we have integrated the ODEs of the TOV system by a Discontinuous Galerkin solver like in [27], obtaining a total mass $M=1.4\,M_{\odot}$ and a radius $R=14.15\,{\textrm{km}}$. A small perturbation in the pressure is then added to the initial model, which is evolved with the well–balanced version of the CWENO5 scheme. The logical computational domain is given by the box $\Omega_{\ell}=[-20;20]^{3}$, covered by a uniform grid formed by $80\times 80\times 80$ points. This is also the first test where we have used the re-mapping algorithm for the numerical grid described in Sect. III.2, producing a physical computational domain that extends up to $\sim 300M$ in each direction. In particular, the parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.28$, $b=-8.4$, $c=85.0$, $d=-280.0$, $\xi_{c}=10$.
We also stress that in this test we take full advantage of the new algorithm proposed by [27], which allows to convert from the conserved to the primitive variables even in a $\rho=0$ atmosphere in the exterior of the star.
Fig. 8 reports the results of this set–up by showing the one dimensional profiles of various representative quantities at the final time $t=1000$, compared to the exact solution. The small deviations of the velocity and of the scalar $K$ from the equilibrium values are precisely due to the initial perturbation introduced. Fig. 9, on the other hand, shows the evolution of the Einstein constraints and of the central density.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: Solution for the TOV star obtained with the Well-Balanced CWENO5 scheme.
Refer to caption Refer to caption
Figure 9: Time evolution of the Einstein constraints (left panel) and of the central density (right panel) for the 3D TOV star.
IV.7 Single puncture black hole
As next test case we consider a single puncture black hole, as originally proposed in [105] (the so–called puncture (or trumpet) solution). The initial condition for the conformal metric tensor is simply the identity matrix, i.e. $\tilde{\gamma}_{ij}=\delta_{ij}=\mathbf{I}$, while the conformal factor at the initial time is set to $\Psi=1+M/(2r)$ with $r=\left\|\mathbf{x}-\mathbf{x}_{c}\right\|$. The black hole is centered in the origin $\mathbf{x}_{c}=(0,0,0)$, has unit mass $M=1$ and zero spin. The initial extrinsic curvature and the initial shift are set to zero, i.e. $K_{ij}=0$, $\beta^{i}=0$, while the lapse is initialized with $\alpha=\psi^{-2}$. The (logical) three dimensional computational domain is given by $\Omega_{\ell}=[-13,13]^{3}$ covered by $132^{3}$ gridpoints with grid-stretching activated. The parameters of the polynomial stretching in the grid function (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-176,176]^{3}$. To avoid the singularity in $r=0$, the computational mesh is built in such a way that no grid point coincides with the origin $\mathbf{x}_{c}$, i.e. the grid points are located in the centers of the equidistant control volumes of the logical grid described above, see also Section III.1 for details. The gamma–driver is turned on, i.e. $s=1$, and the damping parameter in the gamma driver is set to $\eta=2$. We furthermore set $\mu=0$. We have solved this problem with the CWENO7 scheme using exponent $r=4$ in Eq. (55), $\lambda_{0}=10^{10}$, $\lambda_{C}=10^{6}$, and $\lambda_{L}=\lambda_{R}=1$ in Eq. (51) and stopping the simulation at $t=1000$. The initial condition is imposed as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
Fig. 10 provides a first glance of the results, by reporting the contour plots of the lapse in the $xy$ plane (left panel) as well as the time evolution of the Einstein constraints (right panel). Limited to this test, we have taken the opportunity to perform a close comparison among the results obtained with our new first–order formulation (FO-BSSNOK) and the much more popular second–order formulation (SO-BSSNOK). Hence, Fig. 11 presents such a comparison by reporting the 1D profiles of $\alpha$, $\beta^{1}$, $K$ and $\psi$ in the two cases. Although the numerical schemes adopted are different, namely a seventh order CWENO for FO-BSSNOK and a sixth order linear central finite difference for SO-BSSNOK, the various profiles match perfectly well at the final time $t=1000$. Moreover, also the Einstein constraints, compared in the right panel of Fig. 10, manifest essentially the same trend.
Refer to caption Refer to caption
Figure 10: Solution of the single puncture black hole in a 3D calculation at $t=1000$. Left: contour-plots of the lapse over the $xy$ plane. Right: time evolution of the Einstein constraints.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Comparison among the numerical solutions at time $t=1000$ for the single puncture black hole obtained with (i) our new seventh order CWENO finite difference scheme in the first order FO-BSSNOK system and (ii) a classical sixth order central finite difference scheme applied to the standard second order formulation of BSSNOK (SO-BSSNOK). 1D cuts along the $x_{1}$ axis are shown: lapse $\alpha$ (top left), shift vector component $\beta^{1}$ (top right), trace of the extrinsic curvature $K$ (bottom left) and conformal factor $\psi$ (bottom right).
IV.8 Head on collision of two puncture black holes
Here and in the following Section, we use the procedure initially proposed by [105, 106] to create the initial conditions corresponding to two puncture black holes without excision. The initial metric and the lapse are provided by the TWOPUNCTURES initial data code of [106]. As a first set up, we consider two non–spinning black holes, at rest with respect to each other, and at an initial distance $d=2\,M$. This is the so–called head–on collision. The gamma–driver is necessarily turned on, i.e. $s=1$, again with damping parameter $\eta=2$ and $\mu=0$. We have evolved this configuration with the CWENO7 scheme, over a logical computational domain $\Omega_{\ell}=[-15,15]^{3}$ covered by $132^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-335,335]^{3}$. Also in this test case the initial condition is imposed as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
We have stopped the evolution at $t=1000$, reporting the results in Fig. 12, where we show the contour plots of the scalar $\psi=e^{\phi}$ over the $z=0$ plane. The two black holes merge as expected and at time $t=20$ a single black hole is already formed. The corresponding evolution of the Einstein constraints is instead visible in the left panel of Fig. 14. The constraints slightly increase in time, but only in a very mild form.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12: Head on collision of two black holes obtained with CWENO7.
IV.9 Inspiralling merger of two black holes
Numerical investigations of inspiralling black hole mergers started well before the detection of gravitational waves, and a vast literature exists about their behaviour (see, among the other, [93, 107, 108]). Using again the TWOPUNCTURES library of [106], we place the two black holes at a distance $d=4\,M$ among each other, zero individual spins, opposite linear momenta along the $y$ direction, i.e., $p_{1}=(0,0.45,0)$, $p_{2}=(0,-0.45,0)$, and masses of the two black holes $m_{1}=m_{2}=1$. The gamma–driver is again turned on, i.e. $s=1$, with $\eta=2$ and $\mu=0$. In these conditions we have evolved the system with the CWENO7 scheme over a logical computational domain $\Omega_{\ell}=[-15,15]^{3}$ covered by $132^{3}$ gridpoints, and grid-stretching activated. The parameters of the polynomial stretching in the grid function of Eq. (72) are $a=0.32$, $b=-4.8$, $c=25.0$, $d=-40.0$, $\xi_{c}=5$, producing a physical domain of size $\Omega=[-335,335]^{3}$. We use the initial condition as Dirichlet boundary condition on the boundary of the computational domain $\partial\Omega$.
We have stopped the evolution at $t=1000$, reporting the results in Fig. 13, where we show the contour plots of the scalar $\psi=e^{\phi}$ over the $z=0$ plane. The two black holes merge after a brief inspiral phase and at time $t=40$ a single black hole is already formed. The corresponding evolution of the Einstein constraints is instead visible in the right panel of Fig. 14.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13: Inspiralling merger of two black holes obtained with CWENO7.
Refer to caption Refer to caption
Figure 14: Time evolution of the Einstein constraints. Left: Head on collision of Fig. 12. Right: Inspiralling merger of Fig. 13
V Conclusions
The second–order BSSNOK formulation of the Einstein equations represents a sort of gold standard in NR, but it also contains an unpleasant feature, in so far it requires quite different numerical schemes for the solution of the Einstein sector and of the matter subsystem. More specifically, writing the Einstein equations as a second–order PDE system prevents from applying to them the best numerical schemes developed over the last decades for hyperbolic systems of conservation laws, which have been developed having first–order systems in mind. In this paper we propose two fundamental advancements in this respect:
  • We have provided a new first–order BSSNOK formulation of the Einstein–Euler equations which is non–conservative for the Einstein equations, while of course conservative for the matter equations. Moreover, the full PDE system is provably strongly hyperbolic.
  • We have presented a new high order (verified up to the seventh) path-conservative central WENO finite difference scheme that is able to account for the mixed nature of the first–order PDE system.
Concerning the new first–order formulation, our work has been inspired by [23], from which, however, we differ in the crucial aspect that we carefully avoided to insert first order constraints in the equations. Such terms, we believe, introduce artificial Jordan blocks in the overall PDE system and thus may lead to a loss of strong hyperbolicity.
Concerning the numerical scheme, our work resumes a sequence of works by [59, 60, 61], who proposed central WENO schemes more than twenty years ago for classical gas dynamics. Much more recently, on the other hand, [64] considered a new class of alternative finite difference WENO schemes for the non–conservative Z4 formulation of the vacuum Einstein equations, and to their achievements we are also highly indebted.
Our new implementation can successfully solve all the standard tests of numerical relativity [92], including the long term evolution of a black hole binary system. We believe that the NR community can benefit from this work, both regarding the formulation of the Einstein equations, and the newly proposed numerical scheme, which, being a finite difference method, maintains all the advantages with respect to more expensive DG or finite volume schemes, and yet it can reach any desired high order of accuracy.
VI Acknowledgments
This work was financially supported by i) the Italian Ministry of Education, University and Research (MIUR) in the framework of the PRIN 2022 project High order structure-preserving semi-implicit schemes for hyperbolic equations and via the Departments of Excellence Initiative 2018–2027 attributed to DICAM of the University of Trento (grant L. 232/2016) and ii) by the Departments of Excellence Initiative 2022–2027 attributed to the Department of Mathematics of the University of Roma La Sapienza (grant L. 232/2016).
M.D. and G.P. are members of the INdAM GNCS group in Italy.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high–performance computational resources and support.
We would kindly like to thank Ilya Peshkov, Luciano Rezzolla and Federico Guercilena for inspiring discussions. The authors declare that not a single line of this paper has been written with AI. The Fortran source code of the computer software which has generated the results shown in this paper can be obtained under CC BY-NC 4.0 license from the authors after written request.
Appendix A Eigenvalues and eigenvectors of the new FO-BSSNOK system
The eigenstructure of the Einstein–Euler system can be studied by focusing on the Einstein block only, more specifically by setting to zero all the hydrodynamic variables $(D,S_{1},S_{2},S_{3},E)$, whose eigenvalues and eigenvectors are well known. In addition, we assume for simplicity the case in which the gamma–driver is switched off ($s=0$) and the 1+log gauge condition is adopted, hence getting a reduced system composed of 34 variables, given by
$\tilde{\boldsymbol{Q}}^{T}$ $=$ $\Big{(}\tilde{A}_{11},\tilde{A}_{12},\tilde{A}_{13},\tilde{A}_{22},\tilde{A}_{23},\tilde{A}_{33},K,\tilde{\Gamma}^{1},\tilde{\Gamma}^{2},\tilde{\Gamma}^{3},A_{1},A_{2},A_{3},D_{111},D_{112},D_{113},D_{122},D_{123},D_{133},$ (76)
$D_{211},D_{212},D_{213},D_{222},D_{223},D_{233},D_{311},D_{312},D_{313},D_{322},D_{323},D_{333},P_{1},P_{2},P_{3}\Big{)}\,.$
The eigenvalues of the reduced system in $x_{1}$-direction are
$\lambda_{1,\ldots,5}$ $=$ $+\alpha{\psi}^{-2}\sqrt{\tilde{\gamma}^{11}}-\beta^{1}$ (77)
$\lambda_{6,\ldots,10}$ $=$ $-\alpha{\psi}^{-2}\sqrt{\tilde{\gamma}^{11}}-\beta^{1}$ (78)
$\lambda_{11,\ldots,32}$ $=$ $-\beta^{1}\,,$ (79)
$\lambda_{33}$ $=$ $+\sqrt{2\,\alpha}\,\psi^{-2}\sqrt{\tilde{\gamma}^{11}}-\beta^{1}$ (80)
$\lambda_{34}$ $=$ $-\sqrt{2\,\alpha}\,\psi^{-2}\sqrt{\tilde{\gamma}^{11}}-\beta^{1}$ (81)
while the corresponding eigenvectors are
$\boldsymbol{r}_{1}^{T}$ $=$ $\Big{(}-{\frac{\tilde{\gamma}^{33}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,0,0,{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{33}}{\tilde{\gamma}^{11}}},0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (82)
$\boldsymbol{r}_{2}^{T}$ $=$ $\Big{(}-2\,{\frac{\tilde{\gamma}^{23}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,0,{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{23}}{\tilde{\gamma}^{11}}},0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (83)
$\boldsymbol{r}_{3}^{T}$ $=$ $\Big{(}-{\frac{\tilde{\gamma}^{22}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{22}}{\tilde{\gamma}^{11}}},0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (84)
$\boldsymbol{r}_{4}^{T}$ $=$ $\Big{(}-2\,{\frac{\tilde{\gamma}^{13}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (85)
$\boldsymbol{r}_{5}^{T}$ $=$ $\Big{(}-2\,{\frac{\tilde{\gamma}^{12}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{12}}{\tilde{\gamma}^{11}}},1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (86)
$\boldsymbol{r}_{6}^{T}$ $=$ $\Big{(}2\,{\frac{\tilde{\gamma}^{13}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,-{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (87)
$\boldsymbol{r}_{7}^{T}$ $=$ $\Big{(}2\,{\frac{\tilde{\gamma}^{12}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},-{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{12}}{\tilde{\gamma}^{11}}},1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (88)
$\boldsymbol{r}_{8}^{T}$ $=$ $\Big{(}{\frac{\tilde{\gamma}^{33}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,0,0,-{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{33}}{\tilde{\gamma}^{11}}},0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (89)
$\boldsymbol{r}_{9}^{T}$ $=$ $\Big{(}2\,{\frac{\tilde{\gamma}^{23}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,0,-{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,-2\,{\frac{\tilde{\gamma}^{23}}{\tilde{\gamma}^{11}}},0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (90)
$\boldsymbol{r}_{10}^{T}$ $=$ $\Big{(}{\frac{\tilde{\gamma}^{22}}{\sqrt{\tilde{\gamma}^{11}}{\psi}^{2}}},0,0,-{\frac{\sqrt{\tilde{\gamma}^{11}}}{{\psi}^{2}}},0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{22}}{\tilde{\gamma}^{11}}},0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (91)
$\boldsymbol{r}_{11}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0\Big{)}$ (92)
$\boldsymbol{r}_{12}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{11},2\,\tilde{\gamma}^{12},2\,\tilde{\gamma}^{13},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1,0,0\Big{)}$ (93)
$\boldsymbol{r}_{13}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{13},-2\,{\frac{\tilde{\gamma}_{23}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},-2\,{\frac{\Lambda_{1}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1\Big{)}$ (94)
$\boldsymbol{r}_{14}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,{\frac{\tilde{\gamma}_{33}}{\tilde{\gamma}^{11}}},-{\frac{\tilde{\gamma}_{23}}{\tilde{\gamma}^{11}}},-{\frac{\tilde{\gamma}^{12}}{\tilde{\gamma}^{11}}},1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (95)
$\boldsymbol{r}_{15}^{T}$ $=$ $\Big{(}{\frac{\tilde{\gamma}_{11}}{\tilde{\gamma}_{33}}},{\frac{\tilde{\gamma}_{12}}{\tilde{\gamma}_{33}}},{\frac{\tilde{\gamma}_{13}}{\tilde{\gamma}_{33}}},{\frac{\tilde{\gamma}_{22}}{\tilde{\gamma}_{33}}},{\frac{\tilde{\gamma}_{23}}{\tilde{\gamma}_{33}}},1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (96)
$\boldsymbol{r}_{16}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{12},2\,{\frac{(\tilde{\gamma}^{12})^{2}+\tilde{\gamma}_{33}}{\tilde{\gamma}^{11}}},-2\,{\frac{\tilde{\gamma}_{23}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1,0\Big{)}$ (97)
$\boldsymbol{r}_{17}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}_{23}}{\tilde{\gamma}^{11}}},{\frac{\tilde{\gamma}_{22}}{\tilde{\gamma}^{11}}},-{\frac{\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (98)
$\boldsymbol{r}_{18}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,(\tilde{\gamma}^{11})^{2},\tilde{\gamma}^{11}\,\tilde{\gamma}^{12},\tilde{\gamma}^{11}\,\tilde{\gamma}^{13},0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (99)
$\boldsymbol{r}_{19}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{13},2\,{\frac{\Lambda_{2}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-2\,{\frac{\Lambda_{3}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (100)
$\boldsymbol{r}_{20}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12},2\,(\tilde{\gamma}^{12})^{2}+2\,\tilde{\gamma}_{33},2\,{\frac{\Lambda_{2}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (101)
$\boldsymbol{r}_{21}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}_{12}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-\tilde{\gamma}_{33}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}^{11}}{\tilde{\gamma}_{33}}},-{\frac{\Lambda_{5}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-{\frac{\Lambda_{6}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,{\frac{\tilde{\gamma}_{22}}{\tilde{\gamma}_{33}}},{\frac{\tilde{\gamma}_{23}}{\tilde{\gamma}_{33}}},1,$ (102)
[…]
$\boldsymbol{r}_{22}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,{\tilde{\gamma}^{11}}\,{\tilde{\gamma}^{12}},({\tilde{\gamma}^{12}})^{2},{\frac{\Lambda_{7}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (103)
$\boldsymbol{r}_{23}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}\,,-2\,{\frac{\tilde{\gamma}^{12}\,\left(\tilde{\gamma}_{23}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}\right)}{\tilde{\gamma}^{11}}},-2\,{\frac{\Lambda_{8}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,1,$ (104)
[…]
$\boldsymbol{r}_{24}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,(\tilde{\gamma}^{12})^{2},-2\,{\frac{\Lambda_{9}}{\tilde{\gamma}^{11}}},-2\,{\frac{\tilde{\gamma}^{12}\,\left(\tilde{\gamma}_{23}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}\right)}{\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (105)
$\boldsymbol{r}_{25}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{12}}{\tilde{\gamma}^{11}}},0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0,0\Big{)}$ (106)
$\boldsymbol{r}_{26}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{12}}{\tilde{\gamma}^{11}}},0,0,0,0,0,1,0,0,0,0,0,0,0,0,0,0\Big{)}$ (107)
$\boldsymbol{r}_{27}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,-{\frac{\Lambda_{10}}{\tilde{\gamma}_{33}}},-{\frac{\Lambda_{11}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-{\frac{\Lambda_{12}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,-{\frac{\tilde{\gamma}_{12}\,(\tilde{\gamma}_{23})^{2}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}+\tilde{\gamma}_{12}\,\tilde{\gamma}^{11}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},{\frac{\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},$ (108)
[…]
$\boldsymbol{r}_{28}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,2\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13},-2\,{\frac{\Lambda_{13}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-2\,{\frac{\Lambda_{14}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0,0,0,0\Big{)}$ (109)
$\boldsymbol{r}_{29}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,\tilde{\gamma}^{11}\,\tilde{\gamma}^{13},{\frac{\Lambda_{15}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-{\frac{\Lambda_{16}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0,0,0,0,0\Big{)}$ (110)
$\boldsymbol{r}_{30}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,-2\,{\frac{\Lambda_{17}}{\tilde{\gamma}_{33}}},-2\,{\frac{\Lambda_{14}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-2\,{\frac{\Lambda_{19}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0,0,0\Big{)}$ (111)
$\boldsymbol{r}_{31}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,0,-{\frac{\tilde{\gamma}^{13}}{\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,1,0,0,0,0,0\Big{)}$ (112)
$\boldsymbol{r}_{32}^{T}$ $=$ $\Big{(}0,0,0,0,0,0,0,-{\frac{\Lambda_{20}}{\tilde{\gamma}_{33}}},-{\frac{\Lambda_{21}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},-{\frac{\Lambda_{22}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,{\frac{\tilde{\gamma}_{22}\,\tilde{\gamma}^{13}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},{\frac{\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}}{\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}},0,0,0,0,0,0,0,0,0,0,0,0,1,0,0,0\Big{)}$ (113)
$\boldsymbol{r}_{33}^{T}$ $=$ $\Big{(}2\,{\frac{\sqrt{2}\left(\tilde{\gamma}^{12}\,\tilde{\gamma}_{12}+\tilde{\gamma}^{13}\,\tilde{\gamma}_{13}+2\right)}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},-2\,{\frac{\tilde{\gamma}_{12}\,\sqrt{2}\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},2\,{\frac{\sqrt{2}\left(\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}+\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}\right)}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},-2\,{\frac{\sqrt{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{{\alpha}\,\tilde{\gamma}^{11}}}},-2\,{\frac{\sqrt{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},$ (114)
$-2\,{\frac{\sqrt{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},6\,{\frac{\sqrt{2}\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},8\,\tilde{\gamma}^{11},8\,\tilde{\gamma}^{12},8\,\tilde{\gamma}^{13},12\,{\alpha}^{-1},0,0,2\,{\frac{\tilde{\gamma}^{12}\,\tilde{\gamma}_{12}+\tilde{\gamma}^{13}\,\tilde{\gamma}_{13}+2}{\tilde{\gamma}^{11}}},-2\,\tilde{\gamma}_{12},-2\,\tilde{\gamma}_{13},-2\,\tilde{\gamma}_{22},$
$-2\,\tilde{\gamma}_{23},-2\,\tilde{\gamma}_{33},0,0,0,0,0,0,0,0,0,0,0,0,1,0,0\Big{)}$
$\boldsymbol{r}_{34}^{T}$ $=$ $\Big{(}-2\,{\frac{\sqrt{2}\left(\tilde{\gamma}^{12}\,\tilde{\gamma}_{12}+\tilde{\gamma}^{13}\,\tilde{\gamma}_{13}+2\right)}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},2\,{\frac{\tilde{\gamma}_{12}\,\sqrt{2}\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},-2\,{\frac{\sqrt{2}\left(\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}+\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}\right)}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},2\,{\frac{\sqrt{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{{\alpha}\,\tilde{\gamma}^{11}}}},2\,{\frac{\sqrt{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},$ (115)
$2\,{\frac{\sqrt{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},-6\,{\frac{\sqrt{2}\tilde{\gamma}^{11}}{{\psi}^{2}\sqrt{\alpha\,\tilde{\gamma}^{11}}}},8\,\tilde{\gamma}^{11},8\,\tilde{\gamma}^{12},8\,\tilde{\gamma}^{13},12\,(\alpha)^{-1},0,0,2\,{\frac{\tilde{\gamma}^{12}\,\tilde{\gamma}_{12}+\tilde{\gamma}^{13}\,\tilde{\gamma}_{13}+2}{\tilde{\gamma}^{11}}},-2\,\tilde{\gamma}_{12},-2\,\tilde{\gamma}_{13},-2\,\tilde{\gamma}_{22},$
$-2\,\tilde{\gamma}_{23},-2\,\tilde{\gamma}_{33},0,0,0,0,0,0,0,0,0,0,0,0,1,0,0\Big{)}$
where we have defined
$\Lambda_{1}$ $=$ $2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}-(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}-\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}+$ (116)
$\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-(\tilde{\gamma}_{23})^{2}-\tilde{\gamma}^{11}$
$\Lambda_{2}$ $=$ $(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{4}\tilde{\gamma}^{12}+\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}-2\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{23})^{3}(\tilde{\gamma}^{12})^{2}-2\,{\tilde{\gamma}_{23}}^{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$ (117)
$-\tilde{\gamma}_{13}\,(\tilde{\gamma}^{11})^{2}\tilde{\gamma}^{12}-\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,(\tilde{\gamma}^{12})^{2}-\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}$
$\Lambda_{3}$ $=$ $2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{11}-(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{12}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,(\tilde{\gamma}^{11})^{2}$ (118)
$-(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}^{11})^{2}-\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}(\tilde{\gamma}^{12})^{2}-\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{23})^{2}\tilde{\gamma}_{33}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$
$-(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}-(\tilde{\gamma}^{11})^{2}$
$\Lambda_{5}$ $=$ $3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+3\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}+3\,{\tilde{\gamma}_{13}}^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}^{11})^{2}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}$ (119)
$-\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+2\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+2\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}_{12}\,\tilde{\gamma}^{11}\,(\tilde{\gamma}^{12})^{2}-\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}-\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,(\tilde{\gamma}^{13})^{2}$
$-\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}$
$\Lambda_{6}$ $=$ $3\,{\tilde{\gamma}_{12}}^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+3\,(\tilde{\gamma}_{13})^{2}{\tilde{\gamma}_{22}}^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}+3\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}^{11})^{2}-\tilde{\gamma}_{12}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}$ (120)
$-\tilde{\gamma}_{12}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+5\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+5\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{33}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}_{12}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}\,(\tilde{\gamma}^{13})^{2}-\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}$
$\Lambda_{7}$ $=$ $\Lambda_{2}+\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}$ (121)
$\Lambda_{8}$ $=$ $3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{4}-3\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{3}+4\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+3\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{12}-3\,{\tilde{\gamma}_{13}}^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}$ (122)
$-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}^{11})^{2}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}$
$+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}-2\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}-(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}-\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}$
$\Lambda_{9}$ $=$ $3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}_{33}-3\,(\tilde{\gamma}_{13})^{3}(\tilde{\gamma}_{23})^{3}+3\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{33})^{2}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}-(\tilde{\gamma}^{12})^{3}$ (123)
$\Lambda_{10}$ $=$ $3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}-3\,\tilde{\gamma}_{23}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}_{13})^{3}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}^{11}+(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}^{12})^{2}+\tilde{\gamma}_{13}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+\tilde{\gamma}^{12}$ (124)
$\Lambda_{11}$ $=$ $4\,\tilde{\gamma}_{12}\,{\tilde{\gamma}_{13}}^{3}(\tilde{\gamma}_{23})^{3}-4\,(\tilde{\gamma}_{13})^{4}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}+6\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}-(\tilde{\gamma}_{13})^{4}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}-5\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}$ (125)
$-3\,(\tilde{\gamma}_{13})^{3}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-(\tilde{\gamma}_{13})^{3}\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}^{12})^{2}-5\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}^{12})^{3}$
$+\tilde{\gamma}_{13}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}-(\tilde{\gamma}^{12})^{2}$
$\Lambda_{12}$ $=$ $6\,(\tilde{\gamma}_{12})^{2}(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{3}-8\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,(\tilde{\gamma}_{13})^{4}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}_{23}+4\,(\tilde{\gamma}_{12})^{2}(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}+4\,(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}$ (126)
$-2\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}-4\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}-2\,(\tilde{\gamma}_{13})^{3}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}^{12}+2\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}$
$-\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}^{12})^{2}-4\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,(\tilde{\gamma}^{13})^{2}-2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}+\tilde{\gamma}_{12}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}$
$+2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}-2\,\tilde{\gamma}_{13}\,\tilde{\gamma}^{12}\,(\tilde{\gamma}^{13})^{2}-\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$
$\Lambda_{13}$ $=$ $\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}\,(\tilde{\gamma}^{12})^{2}+\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{3}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{11}+\tilde{\gamma}_{23}\,\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}$ (127)
$\Lambda_{14}$ $=$ $3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{4}-3\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{3}+4\,\tilde{\gamma}_{12}\,{\tilde{\gamma}_{13}}^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+3\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{12}-3\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}$ (128)
$-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}^{11})^{2}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-2\,{\tilde{\gamma}_{13}}^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}$
$+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}-2\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{23}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}-(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}$
$\Lambda_{15}$ $=$ $(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{3}\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{4}\tilde{\gamma}^{12}+\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}\tilde{\gamma}_{33}\,\tilde{\gamma}^{13}-2\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{23})^{3}(\tilde{\gamma}^{12})^{2}-2\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}_{33}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$ (129)
$-\tilde{\gamma}_{13}\,(\tilde{\gamma}^{11})^{2}\tilde{\gamma}^{12}-\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}\,(\tilde{\gamma}^{12})^{2}$
$\Lambda_{16}$ $=$ $\Lambda_{3}+(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}+(\tilde{\gamma}^{11})^{2}$ (130)
$\Lambda_{17}$ $=$ $2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}-(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}-\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}+\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$ (131)
$\Lambda_{19}$ $=$ $3\,(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{4}-3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{3}+3\,{\tilde{\gamma}_{12}}^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{11}-3\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}+(\tilde{\gamma}_{13})^{3}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}^{11}$ (132)
$+(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}-2\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}+2\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}(\tilde{\gamma}^{13})^{2}-\tilde{\gamma}_{12}\,(\tilde{\gamma}_{23})^{3}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}$
$+\tilde{\gamma}_{23}\,\tilde{\gamma}^{12}\,(\tilde{\gamma}^{13})^{2}-\tilde{\gamma}_{12}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}$
$\Lambda_{20}$ $=$ $3\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{12})^{2}-3\,(\tilde{\gamma}_{22})^{2}(\tilde{\gamma}_{13})^{3}+(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}+\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}-5\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-\tilde{\gamma}_{13}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}^{13}$ (133)
$\Lambda_{21}$ $=$ $6\,(\tilde{\gamma}_{12})^{2}{\tilde{\gamma}_{13}}^{2}(\tilde{\gamma}_{23})^{3}-8\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,(\tilde{\gamma}_{13})^{4}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}_{23}+4\,(\tilde{\gamma}_{12})^{2}(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}+4\,(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{12}$ (134)
$-2\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}-4\,\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{23})^{2}\tilde{\gamma}^{13}-2\,{\tilde{\gamma}_{13}}^{3}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}^{12}+2\,(\tilde{\gamma}_{13})^{3}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{12})^{2}\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}\,\tilde{\gamma}^{12}$
$-\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}^{11}\,\tilde{\gamma}^{13}+\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,(\tilde{\gamma}^{12})^{2}-4\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}+(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{23}\,(\tilde{\gamma}^{13})^{2}$
$-2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{2}+\tilde{\gamma}_{12}\,(\tilde{\gamma}^{12})^{2}\tilde{\gamma}^{13}+2\,{\tilde{\gamma}_{13}}^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}_{23}-2\,\tilde{\gamma}_{13}\,\tilde{\gamma}^{12}\,(\tilde{\gamma}^{13})^{2}-\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}^{11}-\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}+\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{13}-\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}$
$\Lambda_{22}$ $=$ $4\,(\tilde{\gamma}_{12})^{3}\tilde{\gamma}_{13}\,(\tilde{\gamma}_{23})^{3}-6\,(\tilde{\gamma}_{12})^{2}(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}_{23})^{2}+2\,{\tilde{\gamma}_{13}}^{4}(\tilde{\gamma}_{22})^{3}+2\,(\tilde{\gamma}_{12})^{3}\tilde{\gamma}_{13}\,\tilde{\gamma}_{23}\,\tilde{\gamma}^{11}-(\tilde{\gamma}_{12})^{2}(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,\tilde{\gamma}^{11}$ (135)
$+\tilde{\gamma}_{12}\,(\tilde{\gamma}_{13})^{2}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}^{12}+(\tilde{\gamma}_{13})^{3}(\tilde{\gamma}_{22})^{2}\tilde{\gamma}^{13}+2\,\tilde{\gamma}_{12}\,\tilde{\gamma}_{13}\,\tilde{\gamma}_{22}\,\tilde{\gamma}^{12}\,\tilde{\gamma}^{13}-3\,(\tilde{\gamma}_{13})^{2}\tilde{\gamma}_{22}\,(\tilde{\gamma}^{13})^{2}+\tilde{\gamma}_{12}\,\tilde{\gamma}^{12}\,(\tilde{\gamma}^{13})^{2}$
$-\tilde{\gamma}_{13}\,(\tilde{\gamma}^{13})^{3}-(\tilde{\gamma}^{13})^{2}$